An athlete performing a long jump leaves the ground at a 27.0° angle and lands 7.80 m away. (a) What was the takeoff speed? (b) If this speed were increased by just 5.0%, how much longer would the jump be?
Answer #1

a)R = [v2sin(2θ)]/g

7.8 = [v2sin(2*27)]/9.8    => v= 9.72m/s

b) v = 9.72+ (9.72*0.05) = 10.206m/s

R=[v2sin(2θ)]/g = [10.2062sin(2*27)]/9.8 = 8.6m